cos^2(47)+cos^2(73)+cos(47)*cos(73)

Grade:11

1 Answers

Rinkoo Gupta
9 years ago
1/2[2cos^(47)+2cos^(73)+2cos(47)cos(73)]
=1/2[1+cos(94)+1+cos(146)+cos(120)+cos(26)]
=1/2[2+cos(94)+cos(146)-(1/2)+cos(26)]
=1/2[3/2+{cos(94)+cos(26)}+cos(146)]
=1/2[3/2+2cos(120)cos(34)+cos(146)]
=1/2[3/2+2(1/2)cos(34)+cos(146)]
=1/2[3/2+cos(34)+cos(146)]
=1/2[3/2+2cos(90)cos(56)]
=1/2[3/2+2(0).cos(56)]
=1/2[3/2+0]
=3/4 Ans.
Thanks & Regards
Rinkoo Gupta
AskIITians Faculty